OpenStudy (anonymous):

HELP Choose the equivalent system of linear equations that will produce the same solution as the one given below. 6x + 2y = -6 3x - 4y = -18

OpenStudy (anonymous):

1) 12x + 4y = -12 15x = -30 2) 8x + 4y = -4 14x = -10 3) 6x - 8y = -36 -6y = -42 4) 6x - y = -15 3y = 9

OpenStudy (anonymous):

@study100

OpenStudy (anonymous):

we can use process of elimination to find x

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

2(6x+2y=-6) 3x-4y=-18 ---------------- now distribute 2 12x+4y=-12 >> add these 2 equations 3x - 4y= -18 ---------------- 15x = -30 >> answer now that you look through abcd, there's only 1 answer with the above equation.

OpenStudy (anonymous):

#1?

OpenStudy (anonymous):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = -1 (1 point) Question 2 options: 1) x = 0 2) x = 2 3) There is no x value as there is no solution. 4) x can be any value as there is an infinite number of solutions. @study100

OpenStudy (anonymous):

yes number 1 I'll look at 2 in a sec

OpenStudy (anonymous):

okay thank you